# An Expanded Massive Multilingual Dataset for High-Performance Language Technologies

### v2

@@ 7 Conclusions and future work
We introduce the HPLT v2 dataset, a large-scale multilingual collection of openly-available monolingual and parallel web-crawled data. We focus on improving the quality of data available for a wide range of languages, and we make our data processing pipelines publicly available for easy reuse. We present extensive data analysis as well as intrinsic and extrinsic evaluation, demonstrating the value of HPLT v2 for various NLP tasks. Further work will focus on expanding language coverage and data quality, particularly for under-served languages, and we plan to release a document-level aligned parallel corpus.
@@
During the evaluation, we discovered that the punctuation for Chinese languages (and probably Korean and Japanese) in HPLT v2 had been normalised incorrectly to its Latin equivalent, causing a drop in measured performance for languages in this script. We will fix this in the next iteration of the HPLT v2 pipeline.
@@
The total amount of hours spent would be roughly 4.4M CPU hours and 106K GPU hours. The most expensive task is the evaluation of our data through generative model training. We mitigate the environmental impact of our work by using one of the most eco-efficient data centres in the world to carry out much of our computation.
@@ Appendix A Comparison of multilingual collections
Dataset Size (TB) Tokens (T) Langs % English Source English only The Pile Gao et al. (2020) 0.8 0.39 1 100 various (c.f Section 2) C4.en Raffel et al. (2020); Dodge et al. (2021) 0.3 0.16 1 100 Common Crawl RefinedWeb Penedo et al. (2023) 2.8 0.6 1 100 Common Crawl Dolma-Web Soldaini et al. (2024) - 2.28 1 100 Common Crawl FineWeb Penedo et al. (2024a) - 15 1 100 Common Crawl Multilingual OSCAR-23.01 Suárez et al. (2019) - 1.1 153 48.43 Common Crawl CC-100 Conneau et al. (2020a) 2.39 0.3 100 18.84 Common Crawl mC4 Xue et al. (2021) - 6.3 101 5.67 Common Crawl ROOTS Laurençon et al. (2022) 1.6 0.4 46 30.03 BigScience Catalogue Data, Common Crawl, OSCAR Glot500-c Imani et al. (2023) 0.6 - 511 *2.16 various (c.f. Appendix C) Serengeti Adebara et al. (2023) 0.042 - 517 - various (c.f. Appendix C) CulturaX Nguyen et al. (2024a) 27 6.3 167 45.13 OSCAR, mC4 MADLAD-400-clean Kudugunta et al. (2024) - 2.6 419 50 Common Crawl MaLA (Ji et al., 2024) - 0.074 939 4 various (c.f. Section 2.1.4) monoHPLT v1.2-dedup de Gibert et al. (2024) 11 5.6 75 41 Common Crawl, Internet Archive HPLT v2 (monolingual, deduplicated) 21 7.6 193 44 Common Crawl, Internet Archive
@@ Appendix E Frequent n-grams
Figures 10 and 11 show the five most frequent $n$-grams (orders 1 to 5) in HPLT v2. In the case of parallel datasets, $n$-grams are selected from the target (non-English) side of the segments. Translation to English is obtained with Google Translate. [^24] https://translate.google.com/
@@ Appendix G Geographic TLDs
Tables 9 and 10 list the most frequent examples of geographic TLDs in the monolingual and parallel HPLT v2 corpora respectively. We make the following observations in addition to those made in the main text:
@@ Appendix H Manual quality inspection
We compared four groups of crawls: among wide IA crawls and CC crawls separately we selected old crawls from 2012-2014 and new crawls from 2017-2020. Among languages spoken by the paper authors, 21 languages were selected for annotation.
@@
Figure 12 shows proportions of unnatural texts for each language and group of crawls. Looking at individual languages, for most of them the group of new CC crawls give a much lower proportion of unnatural texts than other groups. However, since only 50 documents were labelled from each group and language, the confidence intervals are large and statistically significant conclusions cannot be made for each individual language. However, when annotations for all languages are combined (denoted as TOTAL on the figure) it becomes clear that for a random language (among those annotated) a random document has 2x lower probability to be unnatural if it comes from the group of newer CC crawls compared to older CC crawls or any of two groups of IA crawls. For the proportions of pornographic content and documents misclassified by LID we did not observe any consistent differences for different groups of crawls.
@@ I.2 Corpora comparison: Norwegian
Evaluation is performed using NorEval, [^28] https://github.com/ltgoslo/noreval/tree/main an open-source benchmark for Norwegian built upon lm-evaluation-harness Gao et al. (2024). We consider the following ten multiple-choice QA, generative QA, sentence completion, and sentence pair ranking tasks that target different aspects of the model understanding and generation abilities in Norwegian Bokmål and Nynorsk:
  • Commonsense reasoning: performing logical and commonsense reasoning (NorCommonsenseQA, Mikhailov et al., 2025).
  • Norwegian-specific & world knowledge: answering questions about facts and Norwegian culture (NorOpenBookQA and NRK-Quiz-QA, Mikhailov et al., 2025).
  • Norwegian language knowledge: understanding Norwegian punctuation rules (NCB) [^29] https://huggingface.co/datasets/hcfa/ncb and idioms (NorIdiom). [^30] https://huggingface.co/datasets/Sprakbanken/Norwegian_idioms
@@
Following the HPLT v1.2, [^31] https://hplt-project.org/HPLT_D4_1___First_language_models_trained.pdf we use UD treebanks of version 2.13323232https://lindat.mff.cuni.cz/repository/xmlui/handle/11234/1-5287 for most languages, except for Albanian and Georgian. These languages were not used in the HPLT v1.2 report due to missing training and development splits in UD 2.13. However, UD 2.15 does contain the required splits, and we use them. We do not evaluate NER on Maltese, since its WikiAnn training split contains only 100 samples. Table 12 shows detailed MLM evaluation results by language and task.
@@
We compare models trained on HPLT v2, Tatoeba (Tiedemann, 2012, 2020), and the combination of the two datasets. The language selection is the intersection of the languages covered by both datasets. We evaluate the models on the FLORES-200 evaluation benchmark (NLLB Team et al., 2024) using SacreBLEU implementation of BLEU333333nrefs:1|case:mixed|eff:no|smooth:exp|version:2.5.1,
@@
and chrF++343434nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1 metrics (Post, 2018) and COMET-22-DA (Rei et al., 2022).
@@
Tables 13 and 14 present the full results of the MT models for translation into English and from English respectively. For reference, we also include the performance of models trained on the HPLT v1.2 dataset, which shares the same underlying extraction pipeline. Note that we did not perform any language-specific hyper-parameter tuning which possibly led to low scores for a few model instances.
HPLT v2 Tatoeba HPLT v2+Tatoeba HPLT v1.2 BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET eng-afr 39.2 64.5 0.8398 38.3 63.6 0.8398 38.8 63.8 0.8409 eng-azj 12.2 41.0 0.8128 11.3 38.7 0.8074 11.5 38.7 0.8011 eng-bul 38.0 62.4 0.8680 0.9 14.5 0.6774 30.0 51.9 0.8122 eng-ben 16.0 45.2 0.8109 16.6 45.9 0.8282 16.8 46.1 0.8275 eng-cat 37.8 61.0 0.8334 39.8 62.2 0.8440 39.5 62.1 0.8425 38.4 61.7 0.8461 eng-cym 50.4 69.9 0.8611 47.7 67.6 0.8536 48.4 67.8 0.8505 eng-est 24.5 53.8 0.8684 24.5 53.3 0.8599 24.4 53.3 0.8578 23.7 53.4 0.8664 eng-eus 16.5 49.5 0.8215 14.9 47.2 0.8098 14.8 47.1 0.8121 12.1 43.4 0.7674 eng-pes 21.5 47.5 0.7947 23.4 50.0 0.8336 23.6 50.0 0.8349 eng-gle 29.0 53.9 0.7543 30.2 53.9 0.7715 30.8 54.6 0.7717 27.3 52.6 0.7561 eng-glg 30.0 55.7 0.8179 31.4 56.1 0.8302 31.4 56.1 0.8264 27.9 54.0 0.8033 eng-guj 19.3 46.5 0.8066 22.5 49.9 0.8518 22.6 49.9 0.8479 eng-heb 28.1 54.0 0.8320 29.7 55.9 0.8532 29.6 55.4 0.8503 eng-hin 32.0 54.6 0.7612 33.1 55.5 0.7728 32.5 54.9 0.7658 32.8 55.5 0.7621 eng-isl 22.2 47.1 0.7766 22.8 47.1 0.7800 23.1 47.5 0.7859 20.6 45.1 0.7651 eng-jpn 27.0 26.6 0.8244 29.9 30.2 0.8640 29.6 29.9 0.8633 eng-kaz 21.0 51.4 0.8651 16.5 45.1 0.8315 16.9 45.3 0.8347 eng-kan 13.8 43.5 0.7746 19.5 50.8 0.8348 19.2 51.1 0.8369 eng-kor 25.0 31.2 0.8268 26.6 32.2 0.8424 26.5 32.0 0.8402 eng-lvs 26.8 53.1 0.8214 23.9 50.0 0.7898 24.3 50.6 0.7891 eng-mal 0.6 20.2 0.5753 14.4 47.9 0.8438 14.6 48.0 0.8427 eng-mar 11.0 37.9 0.6086 13.9 42.3 0.6808 14.2 42.4 0.6792 eng-zsm 38.3 63.9 0.8580 24.4 52.6 0.8534 25.1 53.2 0.8541 eng-sin 1.2 18.6 0.6289 13.1 41.0 0.8542 13.2 41.2 0.8548 eng-slk 29.3 54.0 0.8280 29.3 53.9 0.8353 29.9 54.5 0.8423 eng-slv 26.8 52.2 0.8295 26.5 52.0 0.8339 27.5 52.6 0.8414 eng-als 27.7 54.2 0.8398 29.9 55.8 0.8659 29.3 55.3 0.8600 27.8 54.6 0.8509 eng-swh 32.5 58.2 0.7965 31.2 57.0 0.8058 31.3 57.0 0.8030 28.4 54.6 0.7743 eng-tel 20.2 51.3 0.8104 22.1 53.7 0.8378 22.7 53.9 0.8383 eng-tha 9.9 40.9 0.7977 8.1 40.6 0.8053 8.7 40.9 0.8053 eng-tur 25.3 53.7 0.8368 27.8 56.4 0.8685 27.5 55.8 0.8638 eng-ukr 26.7 52.6 0.8457 27.2 53.4 0.8532 26.8 52.8 0.8471 eng-urd 18.9 43.2 0.7548 19.3 44.0 0.7537 19.5 44.6 0.7584 eng-uzn 16.3 49.1 0.8397 15.9 47.3 0.8497 17.1 48.8 0.8532 eng-vie 37.8 55.8 0.8358 39.3 57.1 0.8489 38.8 56.6 0.8451
@@
Table 13: MT results for models translating from English, trained on our HPLT v2, Tatoeba (OPUS), a combination of both, and the existing HPLT v1.2 (numbers reported where available).
HPLT v2 Tatoeba HPLT v2+Tatoeba HPLT v1.2 BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET azj-eng 18.5 47.1 0.8290 17.4 44.7 0.8039 18.6 46.2 0.8175 bul-eng 35.5 61.1 0.8556 7.4 32.5 0.5104 34.8 60.6 0.8524 ben-eng 27.9 53.7 0.8468 28.4 53.7 0.8498 29.0 54.2 0.8523 cat-eng 39.2 63.1 0.8478 41.1 64.4 0.8580 40.3 63.9 0.8541 41.0 64.4 0.8676 cym-eng 51.5 70.7 0.8615 50.0 68.8 0.8456 50.6 69.1 0.8455 est-eng 30.3 55.8 0.8517 30.7 56.1 0.8510 30.7 55.6 0.8510 30.6 56.6 0.8611 eus-eng 23.3 49.2 0.8121 22.2 47.5 0.8065 22.0 47.4 0.8042 19.4 45.7 0.7810 pes-eng 31.1 56.4 0.8447 33.7 58.2 0.8585 32.7 57.6 0.8546 gle-eng 34.1 58.8 0.8006 32.3 56.3 0.7754 33.1 57.7 0.7918 29.9 54.9 0.7653 glg-eng 33.7 59.2 0.8374 34.5 59.1 0.8395 35.0 59.9 0.8441 31.4 57.2 0.8236 guj-eng 28.5 54.6 0.8475 32.0 57.0 0.8646 33.0 57.6 0.8667 heb-eng 38.2 62.2 0.8534 39.7 62.9 0.8622 40.4 63.6 0.8665 hin-eng 34.7 59.5 0.8701 35.8 60.1 0.8739 36.9 61.0 0.8773 35.2 59.9 0.8741 isl-eng 29.0 53.4 0.8189 29.0 52.8 0.8163 28.7 52.8 0.8136 25.3 50.0 0.7815 jpn-eng 19.9 46.8 0.8255 24.6 52.5 0.8628 23.6 50.6 0.8533 kaz-eng 27.0 53.4 0.8403 22.6 47.8 0.8003 22.6 47.7 0.7998 kan-eng 3.8 24.5 0.6246 27.9 53.4 0.8391 27.4 53.2 0.8396 kor-eng 24.1 51.3 0.8458 25.7 52.7 0.8586 25.8 52.7 0.8578 lvs-eng 29.3 56.0 0.8368 25.0 50.9 0.7862 26.6 53.3 0.8113 mal-eng 2.9 23.5 0.5978 26.4 51.7 0.8342 26.4 51.9 0.8363 mar-eng 23.8 49.8 0.8063 26.1 51.9 0.8299 26.9 52.2 0.8320 zsm-eng 37.2 61.3 0.8561 38.5 61.8 0.8579 38.0 61.7 0.8583 sin-eng 3.0 24.2 0.5979 26.0 51.2 0.8382 26.9 51.9 0.8418 slk-eng 31.6 58.1 0.8456 32.7 58.6 0.8487 33.2 59.0 0.8535 slv-eng 29.2 55.0 0.8371 28.7 54.4 0.8345 29.7 55.6 0.8402 als-eng 32.1 58.6 0.8453 33.7 58.8 0.8449 34.8 59.8 0.8534 31.7 58.3 0.8468 swh-eng 35.3 57.8 0.8086 34.3 56.3 0.7979 33.5 55.6 0.7932 27.2 51.0 0.7542 tel-eng 30.2 55.3 0.8328 31.5 55.9 0.8438 31.9 56.4 0.8446 tha-eng 24.9 52.3 0.8452 22.9 51.0 0.8382 23.7 51.7 0.8411 tur-eng 29.5 54.9 0.8392 32.2 57.3 0.8622 32.7 57.4 0.8602 ukr-eng 33.1 58.7 0.8444 33.4 59.2 0.8470 33.9 59.6 0.8478 urd-eng 26.3 52.1 0.8138 26.2 50.9 0.8097 27.4 52.0 0.8144 uzn-eng 24.8 51.5 0.8110 23.6 48.6 0.7990 24.8 50.0 0.8064 vie-eng 32.0 56.4 0.8514 33.5 57.9 0.8602 32.9 57.2 0.8543
@@
Table 14: MT results for models translating into English, trained on our HPLT v2, Tatoeba (OPUS), a combination of both, and the existing HPLT v1.2 (numbers reported where available).

### v3

@@ 7 Conclusions and future work
This paper introduced the HPLT v2 dataset, a large-scale multilingual collection of openly-available monolingual and parallel web-crawled data. Building on our previous HPLT v1 effort, we focused on improving the quality of data available for a wider range of languages, and we make our data processing pipelines publicly available for easy reuse. We presented extensive data analysis as well as intrinsic and extrinsic evaluation, demonstrating the value of HPLT v2 for various NLP tasks.
Further work will focus on expanding language coverage and data quality, particularly for under-served languages. We also plan to conduct more experiments on the document-level aligned parallel corpus.
@@
During the evaluation, we discovered that the punctuation for Chinese languages (and probably Korean and Japanese) in HPLT v2 had been normalised incorrectly to its Latin equivalent, causing a drop in measured performance for languages in this script. We will fix this in the next iteration of the HPLT pipeline.
@@
The total amount of hours spent would be roughly 4.4M CPU hours and 106K GPU hours. The most expensive task is the evaluation of our data through generative model training. We mitigate the environmental impact of our work by using one of the most eco-efficient HPC clusters in the world (LUMI) to carry out much of our computation. The LUMI supercomputer uses renewable, carbon-neutral energy. [^29] https://lumi-supercomputer.eu/sustainable-future/
Acknowledgements
We acknowledge Tsz Kin Lam, Otto Tarkka, and Maryam Teimouribadelehdareh for their annotation work; we thank Jelmer van der Linde for his help with Go.
This project has received funding from the European Union’s Horizon Europe research and innovation programme under Grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546]. The contents of this publication are the sole responsibility of its authors and do not necessarily reflect the opinion of the European Union.
The authors wish to thank CSC - IT Center for Science, Finland for computational resources and support. The project has also been supported by the Czech MEYS project No. CZ.02.01.01/00/23_025/0008691 and Research Infrastructure project LM2023062.
@@ Appendix A Comparison of multilingual collections
Dataset Size (TB) Tokens (T) Langs % English Source English only The Pile Gao et al. (2020) 0.8 0.39 1 100 various (c.f Section 2) C4.en Raffel et al. (2020); Dodge et al. (2021) 0.3 0.16 1 100 Common Crawl RefinedWeb Penedo et al. (2023) 2.8 0.6 1 100 Common Crawl Dolma-Web Soldaini et al. (2024) - 2.28 1 100 Common Crawl FineWeb Penedo et al. (2024a) - 15 1 100 Common Crawl Multilingual OSCAR-23.01 Suárez et al. (2019) - 1.1 153 48.43 Common Crawl CC-100 Conneau et al. (2020) 2.39 0.3 100 18.84 Common Crawl mC4 Xue et al. (2021) - 6.3 101 5.67 Common Crawl ROOTS Laurençon et al. (2022) 1.6 0.4 46 30.03 BigScience Catalogue Data, Common Crawl, OSCAR Glot500-c Imani et al. (2023) 0.6 - 511 *2.16 various (c.f. Appendix C) Serengeti Adebara et al. (2023) 0.042 - 517 - various (c.f. Appendix C) CulturaX Nguyen et al. (2024a) 27 6.3 167 45.13 OSCAR, mC4 MADLAD-400-clean Kudugunta et al. (2024) - 2.6 419 50 Common Crawl MaLA (Ji et al., 2024) - 0.074 939 4 various (c.f. Section 2.1.4) monoHPLT v1.2-dedup de Gibert et al. (2024) 11 5.6 75 41 Common Crawl, Internet Archive HPLT v2 (monolingual, deduplicated) 21 7.6 193 44 Common Crawl, Internet Archive
@@ Appendix E Frequent n-grams
Figures 9 and 10 show the 5 most frequent $n$-grams (orders 1 to 5) in HPLT v2. In the case of parallel datasets, $n$-grams are selected from the target (non-English) side of the segments. Translation to English is obtained with Google Translate. [^30] https://translate.google.com/
@@ Appendix G Geographic TLDs
Tables 10 and 11 list the most frequent examples of geographic TLDs in the monolingual and parallel HPLT v2 corpora respectively. We make the following observations in addition to those made in the main text:
@@ Appendix H Manual quality inspection
We compared four groups of crawls: among wide IA crawls and CC crawls separately we selected old crawls from 2012-2014 and new crawls from 2017-2020. Among languages spoken by the paper authors, 22 languages were selected for annotation.
@@
Figure 11 shows proportions of unnatural texts for each language and group of crawls. Looking at individual languages, for most of them the group of new CC crawls give a much lower proportion of unnatural texts than other groups. However, since only 50 documents were labelled from each group and language, the confidence intervals are large and statistically significant conclusions cannot be made for each individual language. However, when annotations for all languages are combined (denoted as TOTAL on the figure) it becomes clear that for a random language (among those annotated) a random document has about 2x lower probability to be unnatural if it comes from the group of newer CC crawls compared to older CC crawls or any of two groups of IA crawls. For the proportions of pornographic content and documents misclassified by LID we did not observe any consistent differences for different groups of crawls.
@@ I.2 Corpora comparison: Norwegian
Evaluation is performed using NorEval Mikhailov et al. (2025a), an open-source benchmark for Norwegian built upon LM Evaluation Harness Gao et al. (2024). We consider the following ten multiple-choice QA, generative QA, sentence completion, and sentence pair ranking tasks that target different aspects of the model understanding and generation abilities in Norwegian Bokmål and Nynorsk:
  • Commonsense reasoning: performing logical and commonsense reasoning (NorCommonsenseQA, Mikhailov et al., 2025b).
  • Norwegian-specific & world knowledge: answering questions about facts and Norwegian culture (NorOpenBookQA and NRK-Quiz-QA, Mikhailov et al., 2025b).
  • Norwegian language knowledge: understanding Norwegian punctuation rules (NCB, Mikhailov et al., 2025a) [^34] https://huggingface.co/datasets/hcfa/ncb and idioms (NorIdiom, Mikhailov et al., 2025a). [^35] https://huggingface.co/datasets/Sprakbanken/Norwegian_idioms
@@
Following the HPLT v1.2, [^36] https://hplt-project.org/HPLT_D4_1___First_language_models_trained.pdf we use UD treebanks of version 2.13373737https://lindat.mff.cuni.cz/repository/xmlui/handle/11234/1-5287 for most languages, except for Albanian and Georgian. These languages were not used in the HPLT v1.2 report due to missing training and development splits in UD 2.13. However, UD 2.15 does contain the required splits, and we use them. We do not evaluate NER on Maltese, since its WikiAnn training split contains only 100 samples. Table 13 shows detailed MLM evaluation results by language and task.
@@
We compare models trained on HPLT v2, Tatoeba (Tiedemann, 2012, 2020), and the combination of the two datasets. The language selection is the intersection of the languages covered by both datasets. We evaluate the models on the FLORES-200 evaluation benchmark (NLLB Team et al., 2024) using SacreBLEU implementation of BLEU383838nrefs:1|case:mixed|eff:no|smooth:exp|version:2.5.1,
@@
and chrF++393939nrefs:1|case:mixed|eff:yes|nc:6|nw:0|space:no|version:2.5.1 metrics (Post, 2018) and COMET-22-DA (Rei et al., 2022).
@@
Tables 14 and 15 present the full results of the MT models for translation into English and from English respectively. For reference, we also include the performance of models trained on the HPLT v1.2 dataset, which shares the same underlying extraction pipeline. Note that we did not perform any language-specific hyper-parameter tuning which possibly led to low scores for a few model instances.
HPLT v1.2 HPLT v2 Tatoeba (OPUS) HPLT v2+OPUS BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET en-af 39.2 64.5 0.8398 38.3 63.6 0.8397 38.8 63.8 0.8409 en-ar 26.6 55.0 0.8442 en-az 12.2 41.0 0.8128 11.3 38.7 0.8074 11.5 38.7 0.8011 en-be 11.6 39.0 0.7756 11.2 37.4 0.7767 11.8 38.2 0.7800 en-bg 38.0 62.4 0.8680 0.9 14.5 0.6774 30.0 51.9 0.8122 en-bn 16.0 45.2 0.8109 16.6 45.9 0.8282 16.8 46.1 0.8275 en-bs 4.7 26.0 0.4314 26.8 53.9 0.8672 en-ca 38.4 61.7 0.8461 37.8 61.0 0.8334 39.8 62.2 0.8440 39.5 62.1 0.8425 en-cy 50.4 69.9 0.8611 47.7 67.6 0.8536 48.4 67.8 0.8506 en-eo 27.2 54.7 0.8264 25.5 54.1 0.8500 25.9 54.4 0.8523 en-et 23.7 53.4 0.8664 24.5 53.8 0.8684 24.5 53.3 0.8600 24.4 53.3 0.8578 en-eu 12.1 43.4 0.7674 16.5 49.5 0.8215 14.9 47.2 0.8098 14.8 47.1 0.8122 en-fa 21.5 47.5 0.7947 23.4 50.0 0.8336 23.6 50.0 0.8349 en-fi 21.3 51.1 0.8709 21.3 51.2 0.8725 22.1 51.6 0.8752 en-ga 27.3 52.6 0.7561 29.0 53.9 0.7543 30.2 53.9 0.7715 30.8 54.6 0.7717 en-gl 27.9 54.0 0.8033 30.0 55.7 0.8179 31.4 56.1 0.8302 31.4 56.1 0.8264 en-gu 19.3 46.5 0.8066 22.5 49.9 0.8518 22.6 49.9 0.8479 en-he 28.1 54.0 0.8320 29.7 55.9 0.8532 29.6 55.4 0.8503 en-hi 32.8 55.5 0.7621 32.0 54.6 0.7612 33.1 55.5 0.7728 32.5 54.9 0.7658 en-hr 27.5 53.7 0.8504 en-is 20.6 45.1 0.7651 22.2 47.1 0.7766 22.8 47.1 0.7800 23.1 47.5 0.7859 en-ja 27.0 26.6 0.8244 29.9 30.2 0.8640 29.6 29.9 0.8633 en-kk 21.0 51.4 0.8651 16.5 45.1 0.8315 16.9 45.3 0.8347 en-kn 13.8 43.5 0.7746 19.5 50.8 0.8348 19.2 51.1 0.8369 en-ko 25.0 31.2 0.8268 26.6 32.2 0.8424 26.5 32.0 0.8402 en-lt 22.9 50.9 0.8214 25.0 52.7 0.8486 24.4 52.5 0.8417 en-lv 26.8 53.1 0.8214 23.9 50.0 0.7898 24.3 50.6 0.7891 en-mk 32.2 58.6 0.8478 33.9 60.1 0.8718 34.1 60.1 0.8660 en-ml 0.6 20.2 0.5753 14.4 47.9 0.8438 14.6 48.0 0.8427 en-mr 11.0 37.9 0.6086 13.9 42.3 0.6808 14.2 42.4 0.6792 en-ms 38.3 63.9 0.8580 24.4 52.6 0.8534 25.1 53.2 0.8540 en-mt 34.7 63.2 0.7072 35.8 64.1 0.7102 36.0 64.4 0.7111 en-nb 33.0 58.2 0.8661 en-ne 0.9 20.9 0.4880 12.9 42.8 0.7404 12.8 43.1 0.7386 en-nn 22.8 47.1 0.7766 en-si 1.2 18.6 0.6289 13.1 41.0 0.8542 13.2 41.2 0.8548 en-sk 29.3 54.0 0.8279 29.3 53.9 0.8353 29.9 54.5 0.8423 en-sl 26.8 52.2 0.8295 26.5 52.0 0.8339 27.5 52.6 0.8414 en-sq 27.8 54.6 0.8509 27.7 54.2 0.8398 29.9 55.8 0.8659 29.3 55.3 0.8600 en-sr 32.2 57.5 0.8512 en-sw 28.4 54.6 0.7743 32.5 58.2 0.7964 31.2 57.0 0.8058 31.3 57.0 0.8031 en-ta 14.1 46.7 0.8139 16.7 50.3 0.8565 16.4 50.1 0.8553 en-te 20.2 51.3 0.8104 22.1 53.7 0.8378 22.7 53.9 0.8383 en-th 9.9 40.9 0.7977 8.1 40.6 0.8053 8.7 40.9 0.8053 en-tr 25.3 53.7 0.8368 27.8 56.4 0.8685 27.5 55.8 0.8638 en-uk 26.7 52.6 0.8457 27.2 53.4 0.8532 26.8 52.8 0.8471 en-ur 18.9 43.2 0.7548 19.3 44.0 0.7537 19.5 44.6 0.7584 en-uz 16.3 49.1 0.8397 15.9 47.3 0.8497 17.1 48.8 0.8532 en-vi 37.8 55.8 0.8358 39.3 57.1 0.8489 38.8 56.6 0.8451 en-xh 12.0 44.2 0.7323 0.0 3.5 0.2317 0.0 3.5 0.1647
@@
Table 14: MT results (BLEU, chrf++, COMET22-DA) for models translating from English, trained on our HPLT v2, Tatoeba (OPUS), a combination of both, and the existing HPLT v1.2 (numbers reported where available).
HPLT v1.2 HPLT v2 Tatoeba (OPUS) HPLT v2+OPUS BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET BLEU chrF++ COMET az-en 18.5 47.1 0.8290 17.4 44.7 0.8039 18.6 46.2 0.8175 be-en 16.1 46.7 0.7886 14.8 44.7 0.7621 15.5 45.8 0.7743 bg-en 35.5 61.1 0.8556 7.4 32.5 0.5104 34.8 60.6 0.8524 bn-en 27.9 53.7 0.8468 28.4 53.7 0.8498 29.0 54.2 0.8523 bs-en 12.8 38.0 0.5882 35.2 59.9 0.8508 ca-en 41.0 64.4 0.8676 39.2 63.1 0.8478 41.1 64.4 0.8580 40.3 63.9 0.8541 cy-en 51.5 70.7 0.8615 50.0 68.8 0.8456 50.6 69.1 0.8455 eo-en 35.5 59.2 0.8362 35.6 59.2 0.8437 35.6 59.3 0.8429 et-en 30.6 56.6 0.8611 30.3 55.8 0.8517 30.7 56.1 0.8510 30.7 55.6 0.8510 eu-en 19.4 45.7 0.7810 23.3 49.2 0.8121 22.2 47.5 0.8064 22.0 47.4 0.8042 fa-en 31.1 56.4 0.8447 33.7 58.2 0.8585 32.7 57.6 0.8546 fi-en 26.6 52.4 0.8442 26.6 51.9 0.8451 26.2 51.6 0.8421 ga-en 29.9 54.9 0.7653 34.1 58.8 0.8006 32.3 56.3 0.7754 33.1 57.7 0.7918 gl-en 31.4 57.2 0.8236 33.7 59.2 0.8374 34.5 59.1 0.8395 35.0 59.9 0.8441 gu-en 28.5 54.6 0.8475 32.0 57.0 0.8646 33.0 57.6 0.8667 he-en 38.2 62.2 0.8534 39.7 62.9 0.8622 40.4 63.6 0.8665 hi-en 35.2 59.9 0.8741 34.7 59.5 0.8701 35.8 60.1 0.8738 36.9 61.0 0.8773 hr-en 31.7 56.4 0.8389 is-en 25.3 50.0 0.7815 29.0 53.4 0.8189 29.0 52.8 0.8163 28.7 52.8 0.8136 ja-en 19.9 46.8 0.8255 24.6 52.5 0.8628 23.6 50.6 0.8533 kk-en 27.0 53.4 0.8403 22.6 47.8 0.8003 22.6 47.7 0.7998 kn-en 3.8 24.5 0.6246 27.9 53.4 0.8391 27.4 53.2 0.8396 ko-en 24.1 51.3 0.8458 25.7 52.7 0.8586 25.8 52.7 0.8578 lt-en 26.5 51.9 0.8138 27.2 52.6 0.8224 27.0 52.2 0.8201 lv-en 29.3 56.0 0.8368 25.0 50.9 0.7862 26.6 53.3 0.8113 mk-en 37.0 61.4 0.8522 38.2 62.0 0.8558 38.8 62.6 0.8596 ml-en 2.9 23.5 0.5978 26.4 51.7 0.8342 26.4 51.9 0.8363 mr-en 23.8 49.8 0.8063 26.1 51.9 0.8299 26.9 52.2 0.8320 ms-en 37.2 61.3 0.8561 38.5 61.8 0.8579 38.0 61.7 0.8583 mt-en 45.0 68.4 0.7777 47.1 68.5 0.7892 47.5 68.9 0.7895 nb-en 37.3 61.0 0.8540 ne-en 22.1 47.9 0.7929 26.6 52.0 0.8408 27.7 52.8 0.8436 nn-en 32.9 55.8 0.7945 si-en 3.0 24.2 0.5979 26.0 51.2 0.8381 26.9 51.9 0.8418 sk-en 31.6 58.1 0.8456 32.7 58.6 0.8486 33.2 59.0 0.8535 sl-en 29.2 55.0 0.8371 28.7 54.4 0.8345 29.7 55.6 0.8402 sq-en 31.7 58.3 0.8468 32.1 58.6 0.8453 33.7 58.8 0.8448 34.8 59.8 0.8534 sr-en 37.4 62.7 0.8544 sw-en 27.2 51.0 0.7542 35.3 57.8 0.8086 34.3 56.3 0.7979 33.5 55.6 0.7932 ta-en 24.0 49.6 0.8068 23.4 49.0 0.8139 24.3 49.5 0.8154 te-en 30.2 55.3 0.8328 31.5 55.9 0.8438 31.9 56.4 0.8446 th-en 24.9 52.3 0.8452 22.9 51.0 0.8381 23.7 51.7 0.8410 tr-en 29.5 54.9 0.8392 32.2 57.3 0.8622 32.7 57.4 0.8602 uk-en 33.1 58.7 0.8444 33.4 59.2 0.8470 33.9 59.6 0.8478 ur-en 26.3 52.1 0.8138 26.2 50.9 0.8097 27.4 52.0 0.8144 uz-en 24.8 51.5 0.8110 23.6 48.6 0.7990 24.8 50.0 0.8064 vi-en 32.0 56.4 0.8514 33.5 57.9 0.8602 32.9 57.2 0.8543 xh-en 22.1 45.5 0.6703 0.1 11.8 0.2871 0.2 13.1 0.2628
@@
Table 15: MT results (BLEU, chrf++, COMET22-DA) for models translating into English, trained on our HPLT v2, Tatoeba (OPUS), a combination of both, and the existing HPLT v1.2 (numbers reported where available).